The concentration of cation vacancies in NaCl crystal doped with CdCl2 is found to be 6.02*10^16 /mol. What is the concentration of CdCl2 added to it.(2 Marks)
Answer #1
Number of cation vacancies=6.02*10 16 mol-1 NaCl=10-7*6.02*10 23mol-1 NaCl=10-7 mole/mole of NaCl. Hence one mole of NaCl is doped with 10-7 mole of CdCl2
source: Topper of Johnstone hr sec school Imphal
answered by: Alexander Ninthoujam